A person travels by car from one city to another with different constant speeds between pairs of cities. She drives for 30.0 min at 65.0 km/h, 5.0 min at 100.0 km/h, and 50.0 min at 30.0 km/h and spends 35.0 min eating lunch and buying gas.
Answer #1

(B)Total distance travelled = =(30/60)*65+(5/60)*100+(50/60)*30=65.83km

Total time =(30+5+50+35)/60=2 hr

(A)Averge speed = Total distance/Total time= 65.83/2 = 32.92 km/h
Answer #2

Here. A and B represents the starting and end point respectively.